A guitar is tuned to play a note at 424 Hz. If the guitar is in a moving car and the note has a perceived frequency of 409 Hz, what is the speed of the car? Assume the car is moving directly away from a listener on the sidewalk. m/s
Answer #1

we know, speed of sound, v = 340 m/s

velocity of source, vs = ?

f = 424 hz (actual frequency)

f' = 409 hz (apparent freuqncy)


we know, f' = f*v/(v + vs)

v + vs = f*v/f'

vs = f*v/f' - v

= 424*340/409 - 340

= 12.47 m/s <<<<<<<<<<<------------Answer